# Hundred-Kilobyte Lookup Tables for Efficient Single-Image Super-Resolution

### v1

@@ Introduction
Performance and Storage In Table 4, we compared our proposed HKLUT with well-recognized SISR approaches such as interpolation, sparse-coding, and CNNs, as well as the latest LUT-based methods. Our approach has a clear advantage in storage, which aligns with the objective of this paper. HKLUT-S, being the smallest parameterized model, only requires 100KB, which is 10×10$\times$10 × less than the second smallest SRLUT (1.27MB) with a much better performance. Both HKLUT-S and HKLUT-L outperform all classical baselines, including A+, which is significantly larger than HKLUT-S (15.17MB vs. 100KB). Although CNN-based approaches exhibit good performance in terms of PSNR and SSIM, they often require larger storage. For example, RRDB is 650×650$\times$650 × larger than our HKLUT-S, making it difficult to deploy on edge devices. With respect to LUT-based approaches, our approaches not only take up the least storage, but they also upkeep comparable performance. HKLUT-L outperforms MuLUT-SDY on the Set5 dataset (30.41dB vs. 30.40dB) with 35×35$\times$35 × less storage. Despite being 70×70$\times$70 × smaller, HKLUT-S outperforms SPLUT-M by a significant margin (30.35dB vs. 30.23dB on Set5).
@@
Energy & Peak Memory. Following the setting in MuLUT (Li et al. 2022), Fig. 6 shows the theoretical energy costs and peak memory usages of FSRCNN (32.70dB) and several LUT-based methods (31.73dBsimilar-$\sim$32.35dB) for generating a $1280\times 720$ image through 2×2$\times$2 × SR. We estimated the theoretical energy cost using the same protocol as AdderSR (Song et al. 2021), and computed the peak memory usage using memray [^1] Measured with memray (https://github.com/bloomberg/memray) on a desktop CPU.. For LUT-based approaches, we used their smallest variants. One of the major advantages of LUT-based methods is their ability to drastically reduce energy and memory consumption. Even FSRCNN, the simplest CNN for SISR, consumes 900×900$\times$900 × and 7.5×7.5$\times$7.5 × more energy and memory compared to HKLUT. The feature map aggregation module of SPLUT not only instantiates larger LUTs but also requires floating-point operations, resulting in increased energy consumption and peak memory usage. Meanwhile, both SRLUT and MuLUT consume a significant amount of energy for the interpolation operation. In contrast, HKLUT gets rid of energy-intensive interpolation and solely relies on integer operations. Fig. 6 shows that HKLUT exhibits the lowest energy cost and peak memory usage amongst all competitors.
Figure 7: Qualitative results on selected images from five benchmark datasets.
Runtime. Using the same settings as for calculating energy and peak memory, we compared the runtimes of FSRCNN and LUT-based approaches on both a desktop CPU and the Raspberry Pi 4 Model B. To ensure a fair comparison, we used the official Python code released by the authors (Jo and Kim 2021; Li et al. 2022; Ma et al. 2022), without relying on native mobile implementation or any speedup tricks. We recorded the average runtime of generating a $1280\times 720$ image from a $640\times 360$ image in Table 5. It is clear that the 4-simplex interpolation bottlenecks the inference speed of SRLUT and MuLUT at the final stage. SPLUT and our approaches, featuring parallel branches, exhibit shorter inference times on Raspberry Pi and CPU. Unsurprisingly, HKLUT achieves the fastest inference speed.
@@
Qualitative Results
Fig. 7 presents a visual comparison between the bicubic interpolation baseline and different LUT-based approaches on images selected from five benchmark datasets. The results clearly demonstrate that LUT-based approaches improve sharpness compared to bicubic interpolation which generates blurry images. The 5th row of images (Manga109) shows that the SRLUT approach exhibits severe rippling artifacts around the edges, which may be caused by its limited RF. HKLUT-S produces similar sharpness and clarity to MuLUT-SDY-X2 and SPLUT-L, despite being more than 40×40$\times$40 × and 180×180$\times$180 × smaller, respectively. Although A+ brings fewer artifacts, it is computation-heavy and not able to show clearer textures. Overall, the results show that LUT-based approaches outperform bicubic interpolation qualitatively, and our proposed HKLUT has been proven to be a highly competitive alternative to its larger-size LUT-based counterparts, yet at a drastically reduced storage cost.
Conclusion
This paper targets minimalist LUTs for efficient LUT-based SISR. We propose a new class of models called HKLUTs employing kernels with fewer input pixels while retaining the RF, resulting in an exponential reduction in storage without compromising performance. In particular, HKLUT consists of an asymmetric parallel structure that employs big-little kernel patterns to adapt to the underlying characteristics of MSB and LSB branches. A progressive multistage architecture, without expensive interpolation between stages, is devised to reduce storage by half. By seamlessly fusing these innovations, HKLUT models set the first record to attain $>30$dB on the Set 5 dataset at merely one-tenth of 1MB storage. Extensive experiments verify that HKLUT is a versatile alternative to previous large LUT-based approaches on resource-constrained devices.
@@ Conclusion
Details of Neural Nets during Training
During training, each look-up table (LUT) is replaced with a simple convolutional neural network (CNN), as shown in Fig. 8. The CNN consists of one convolution (CONV) layer with a $1\times K$ kernel, and five $1\times 1$ CONV layers, interleaved with RELU (Glorot, Bordes, and Bengio 2011) nonlinearity. No dense and residual connections are used here. $K$, $C$, and $S$ represent the number of input pixels to the LUT, the output channel number of the CONV layer (i.e., 64), and the upscaling factor, respectively. Since the LUT only accepts a finite number (K) of input pixels, all convolution kernels are of size $1\times 1$, except for the first one, which is of size $1\times K$. If the kernel shape of the first convolution is neither horizontal nor vertical, the pixels are first sampled according to the defined index pattern and then reshaped to $1\times K$. After the last CONV layer, pixel shuffle (Shi et al. 2016b) is applied to generate the high-resolution (HR) image. A Tanh activation function is used at the end to project the output to a range between -1 and 1. To cache CNN outputs into a LUT, they are converted from FP32 to INT8. This is done by multiplying the FP32 values by 127 to ensure they fit within the INT8 range (-128 to 127). The resulting numbers are rounded down to the nearest integer and stored in the LUT as INT8.
@@
Effective Receptive Field
As elaborated in the manuscript, SPLUT (Ma et al. 2022) neglects the disparity between the information contained in the most-significant-bit (MSB) and the least-significant-bit (LSB) branches, despite their identical theoretical receptive fields (e.g., $5\times 5$). We hypothesize that using the same sets of LUTs for both MSB and LSB branches may not be the optimal choice. To find an ideal solution, we visualize the effective receptive fields (ERFs) of both branches and introduce an asymmetric structure based on this observation, aiming to substantially reduce the size of LUTs. In Fig. 2, we provide additional demonstrations of the ERFs of various images across datasets, which confirm that the MSB branch has a larger ERF while the LSB branch maintains a more compact one.
LUT Design Comparison
@@
Architectural Design Comparison
This section compares the architectural designs of different LUT-based approaches for SISR. Fig. 11 shows block diagrams that illustrate the overall architectures of these approaches. In this context, LUT refers to a CNN with a structure similar to that of the CNN shown in Fig. 1. The prefix or subscript describes the kernel shape of the first CONV layer of this CNN.
Fig. 11 (a) depicts the overall structure of SRLUT (Jo and Kim 2021). SRLUT consists of only a single LUT, which corresponds to a CNN with the first layer using a $2\times 2$ square kernel. During both training and inference, rotation ensemble is used to expand the receptive field (RF) from $2\times 2$ to $3\times 3$. MuLUT (Li et al. 2022), as illustrated in Fig. 4 (c), extends SRLUT by using multiple LUTs with different input kernel shapes in parallel. These kernels have complementary indexing patterns, as they can fill up the entire $5\times 5$ area using the rotation ensemble trick. The same structure can be stacked in a sequential manner to form a multistage structure to further enlarge the RF.
Fig. 4 (b) shows the overall structure of SPLUT (Ma et al. 2022). SPLUT first divides an 8-bit image into two branches: one with the most significant 4 bits (MSB) and the other with the least significant 4 bits (LSB) to eliminate interpolation at the final stage. To maintain consistency, we follow the terminology from SPLUT (Ma et al. 2022). Here, $\rm{LUT}_{HW}$ is referred to as spatial lookup, which uses the $2\times 2$ square kernel as the S-LUT in Fig. 4 (a) and (c), but with multi-channel outputs. The number of output channels can be either 4 (SPLUT-S), 8 (SPLUT-M), or 16 (SPLUT-L). The multi-channel feature maps are then grouped along the channel dimension and fed to either vertical or horizontal aggregation. The aggregated features are then fed to $\rm{LUT}_{HC}$ and $\rm{LUT}_{WC}$, both of which have a $2\times 2$ input pattern. The input pattern of $\rm{LUT}_{HC}$ is along both the height and channel dimensions, while that of $\rm{LUT}_{WC}$ is along both the width and channel dimensions. The aggregation modules, $\rm{LUT}_{WC}$ and $\rm{LUT}_{HC}$, can be repeated multiple times in series to further expand the RF. The final super-resolved (SR) image is produced by summing the outputs of the two branches.
Figure 11: Comparison of architectural designs. Residual connections are omitted for brevity. (a) SRLUT (Jo and Kim 2021) (b) SPLUT (Ma et al. 2022) (c) MuLUT (Li et al. 2022) (d) HKLUT (Ours)
@@
The proposed HKLUT architecture is illustrated in Fig. 4 (d). Our method utilizes the two-branch structure of SPLUT to eliminate the interpolation step. However, instead of employing the same set of LUTs in both branches, we propose an asymmetric parallel structure that uses smaller LUTs for the LSB branch to reduce storage space. Within each branch, we use multiple LUTs with the rotation ensemble trick similar to MuLUT. However, LUTs of our design have fewer input pixels, resulting in an exponential reduction in size. Summing the results from the two branches yields the final SR image. The entire structure is treated as a building block. A multistage structure is constructed by repeating the whole module. This allows convenient implementation of different progressive upsample strategies to further enhance the results and reduce storage.
Ablation Studies
Ablation on LUT Cascading
In previous work (Ma et al. 2022), a series-parallel structure is proposed to eliminate the need for interpolation. Such parallel structure divides an 8-bit image into MSB and LSB parts to be handled separately, while the series structure refers to cascading LUTs within individual branches. Communication between two branches is only allowed at the end of the model to generate the final SR image by merging information from two branches. The output activations of each LUT in intermediate stages are clipped between [-8, 7] and quantized into 16 levels. However, we find that this strategy is counter-intuitive and may not be optimal. On the contrary, our way of cascading LUTs allows communication between the two branches during stage transition. The output of each stage is clipped between [0, 255] represented in INT8, which allows the generation of meaningful results even in intermediate stages. Table 6 shows that our approach significantly outperforms the multistage structure in SPLUT (Ma et al. 2022) (30.35dB vs. 29.32dB for Set5), both adopting the $2\times 2$ progressive upsampling technique. Note that using the SPLUT way of cascading LUTs without feature aggregation performs even worse than our single-stage method (cf. 1st row of Table 3 in the manuscript).
Ablation on Residual Connection
For computer vision tasks, residual connection (He et al. 2016) is widely used to facilitate the training of neural networks. Various papers (Lai et al. 2017; Ahn, Kang, and Sohn 2018; Ma et al. 2022) have proven the effectiveness of residual connection to maintain the feature precision and improve the SR performance. By using residual connections, the model focuses on reconstructing the residual information between the ground truth HR image and the upsampled image, which intuitively reduces the difficulty of the underlying SR task. Table 7 compares the results of the three commonly used upsampling methods (nearest, bilinear, and bicubic). The effects of the three methods are indistinguishable. To reduce runtime, we selected the nearest interpolation as our upsampling strategy for residual connections.
Method Set5 Set14 BSDS100
PSNR/SSIM PSNR/SSIM PSNR/SSIM
HKLUT-S 30.35/0.859 27.39/0.748 26.73/0.706
SPLUT 29.32/0.836 26.55/0.724 26.15/0.686

### v2

@@ 4.2 Quantitative Results
Performance and Storage In Table 4, we compared our proposed HKLUT with well-recognized SISR approaches such as interpolation, sparse-coding, and CNNs, as well as the latest LUT-based methods. Our approach has a clear advantage in storage, which aligns with the objective of this paper. HKLUT-S, being the smallest parameterized model, only requires 100KB, which is 10×10$\times$10 × less than the second smallest SRLUT (1.27MB) with a much better performance. Both HKLUT-S and HKLUT-L outperform all classical baselines, including A+, which is significantly larger than HKLUT-S (15.17MB vs. 100KB). Although CNN-based approaches exhibit good performance in terms of PSNR and SSIM, they often require larger storage. For instance, RRDB is 650×650$\times$650 × larger than our HKLUT-S, making it challenging to deploy on edge devices. Compared to LUT-based approaches, our methods not only require less storage but also achieve comparable performance. HKLUT-L outperforms MuLUT-SDY on the Set5 dataset (30.41dB vs. 30.40dB) with 35×35$\times$35 × less storage. Despite being 70×70$\times$70 × smaller, HKLUT-S significantly outperforms SPLUT-M (30.35dB vs. 30.23dB).
@@
Energy & Peak Memory. Following the setting in MuLUT Li et al. (2022), Fig. 6 shows the theoretical energy costs and peak memory usages of FSRCNN (32.70dB) and several LUT-based methods (31.73dBsimilar-$\sim$32.35dB) for generating a $1280\times 720$ image through 2×2$\times$2 × SR. We estimated the theoretical energy cost using the same protocol as AdderSR Song et al. (2021), and computed the peak memory usage using memray [^1] Measured with memray (https://github.com/bloomberg/memray) on a desktop CPU.. For LUT-based approaches, we used their smallest variants. One of the major advantages of LUT-based methods is their ability to drastically reduce energy and memory consumption. Even FSRCNN, the simplest CNN for SISR, consumes 900×900$\times$900 × and 7.5×7.5$\times$7.5 × more energy and memory compared to HKLUT. The feature map aggregation module of SPLUT not only instantiates larger LUTs but also requires floating-point operations, resulting in increased energy consumption and peak memory usage. Meanwhile, both SRLUT and MuLUT consume a significant amount of energy for the interpolation operation. In contrast, HKLUT gets rid of energy-intensive interpolation and solely relies on integer operations. Fig. 6 shows that HKLUT exhibits the lowest energy cost and peak memory usage amongst all competitors.
Figure 7: Qualitative results on selected images from Set5, Urban100 and Manga100 datasets.
Runtime. Using the same settings as for calculating energy and peak memory, we compared the runtimes of FSRCNN and LUT-based approaches on both a desktop CPU and the Raspberry Pi 4 Model B. To ensure a fair comparison, we used the official Python code released by the authors, without relying on native mobile implementation or any speedup tricks. We recorded the average runtime of generating a $1280\times 720$ image from a $640\times 360$ image in Table 5. It is clear that the 4-simplex interpolation bottlenecks the inference speed of SRLUT and MuLUT at the final stage. SPLUT and our approaches, featuring parallel branches, exhibit shorter inference times on Raspberry Pi and CPU. Unsurprisingly, HKLUT achieves the fastest inference speed.
@@
4.3 Qualitative Results
Fig. 7 presents a visual comparison between the bicubic interpolation baseline and different LUT-based approaches on images selected from five benchmark datasets. The results clearly demonstrate that LUT-based approaches improve sharpness compared to bicubic interpolation which generates blurry images. The 3rd row (Manga109) shows that the SRLUT approach exhibits severe rippling artifacts around the edges, which may be caused by its limited RF. HKLUT-S produces similar sharpness to MuLUT-SDY-X2 and SPLUT-L, despite being more than 40×40$\times$40 × and 180×180$\times$180 × smaller, respectively.
5 Ablation Studies
5.1 Ablation on Kernel Shapes
To obtain the HDB kernel for our MSB branch, we conducted an ablation study on kernel shape. Instead of exhausting all possible combinations, we employed softmax during training with an annealing temperature to select the most important pixel per column. After training, we evaluated the binarized kernels. We performed 20 runs with different random seeds using a one-stage model, with the LSB branch fixed as the HD kernel. Each run consisted of 20k (10%) training iterations. Next, we selected the top 3 best-performing candidates with non-overlapping kernels and retrained them from scratch for 200k iterations. Fig. 8 shows the 3 best-performing for MSB branch, where HDB (i.e. (a)) ranks as the best.
Figure 8: The 3 best-performing kernel designs for MSB are reported, along with the PSNR/SSIM for 4×4$\times$4 × SR on the Set5 dataset.
Cascading Approach Set5 Set14 BSDS100
PSNR/SSIM PSNR/SSIM PSNR/SSIM
HKLUT 30.35/0.859 27.39/0.748 26.73/0.706
SPLUT 29.32/0.836 26.55/0.724 26.15/0.686
5.2 Ablation on LUT Cascading
In previous work Ma et al. (2022), a series-parallel structure was proposed to remove the need for interpolation. This structure divides an 8-bit image into MSB and LSB parts and processes them separately. The series structure involves cascading LUTs within each branch. Communication between branches only occurs at the end of the model to generate the final SR image by combining information. The output activations of each LUT in intermediate stages are clipped between [-8, 7] and quantized into 16 levels. However, we believe that this approach is counter-intuitive and suboptimal. On the other hand, our method allows communication between branches during stage transition. The output of each stage is clipped between [0, 255] represented in INT8, enabling meaningful results even in intermediate stages. Table 6 demonstrates the superior performance of our approach compared to the multistage structure in SPLUT Ma et al. (2022) (30.35dB vs. 29.32dB for Set5), both using the $2\times 2$ progressive upsampling technique.
6 Conclusion
This paper aims to develop minimalist LUTs for efficient LUT-based SISR. We propose a new class of models called HKLUTs employing kernels with fewer input pixels while retaining the RF, resulting in an exponential reduction in storage without compromising performance. In particular, HKLUT consists of an asymmetric parallel structure that employs big-little kernel patterns to adapt to the underlying characteristics of MSB and LSB branches. A progressive multistage architecture, without expensive interpolation between stages, is devised to reduce storage by half. By seamlessly fusing these innovations, HKLUT models set the first record to attain $>30$dB on the Set 5 dataset at merely one-tenth of 1MB storage. Extensive experiments verify that HKLUT is a versatile alternative to previous large LUT-based approaches on resource-constrained devices.
@@ 6 Conclusion
Appendix A Details of Neural Nets during Training
During training, each look-up table (LUT) is replaced with a simple convolutional neural network (CNN), as shown in Fig. 9. The CNN consists of one convolution (CONV) layer with a $1\times K$ kernel, and five $1\times 1$ CONV layers, interleaved with RELU Glorot et al. [2011] nonlinearity. No dense and residual connections are used here. $K$, $C$, and $S$ represent the number of input pixels to the LUT, the output channel number of the CONV layer (i.e., 64), and the upscaling factor, respectively. Since the LUT only accepts a finite number (K) of input pixels, all convolution kernels are of size $1\times 1$, except for the first one, which is of size $1\times K$. If the kernel shape of the first convolution is neither horizontal nor vertical, the pixels are first sampled according to the defined index pattern and then reshaped to $1\times K$. After the last CONV layer, pixel shuffle Shi et al. [2016b] is applied to generate the high-resolution (HR) image. A Tanh activation function is used at the end to project the output to a range between -1 and 1. To cache CNN outputs into a LUT, they are converted from FP32 to INT8. This is done by multiplying the FP32 values by 127 to ensure they fit within the INT8 range (-128 to 127). The resulting numbers are rounded down to the nearest integer and stored in the LUT as INT8.
@@ Appendix A Details of Neural Nets during Training
Appendix B Effective Receptive Field
As elaborated in the manuscript, SPLUT Ma et al. [2022] neglects the disparity between the information from the most-significant-bit (MSB) and the least-significant-bit (LSB) branches, despite their identical theoretical receptive fields (e.g., $5\times 5$). We hypothesize that using the same set of LUTs for both MSB and LSB branches may not be the optimal choice. To find an ideal solution, we visualize the effective receptive fields (ERFs) of both branches and introduce an asymmetric structure based on this observation, aiming to substantially reduce the size of LUTs. In Fig. 10, we provide additional demonstrations of the ERFs of various images across datasets, confirming that the MSB branch has a larger ERF while the LSB branch maintains a more compact one.
Appendix C LUT Design Comparison
@@
Appendix D Architectural Design Comparison
@@
This section explores various LUT-based SISR architectures. Fig. 12 depicts these architectures. LUT here refers to a CNN similar to that in Fig. 9 with the prefix or subscript describing the kernel shape of the first CONV layer. Fig. 12 (a) depicts the overall structure of SRLUT Jo and Kim [2021], which consists of only a single LUT, corresponding to a CNN with the first layer using a $2\times 2$ square kernel. During training and inference, rotation ensemble is used to expand the receptive field (RF) from $2\times 2$ to $3\times 3$. MuLUT Li et al. [2022], as shown in Fig. 12 (c), extends SRLUT by using multiple LUTs with different input kernel shapes. These kernels have complementary indexing patterns, as they can fill up the entire $5\times 5$ area using the rotation ensemble trick. The same structure can be stacked sequentially to form a multistage structure to further enlarge the RF. Fig. 12 (b) demonstrates SPLUT Ma et al. [2022] which divides an 8-bit image into two branches: one with the most significant 4 bits (MSB) and the other with the least significant 4 bits (LSB) to eliminate interpolation at the final stage. To maintain consistency, we follow the terminology from SPLUT Ma et al. [2022]. Here, $\rm{LUT}_{HW}$ is referred to as spatial lookup, which uses the $2\times 2$ square kernel as the S-LUT in Fig. 12 (a) and (c), but with multi-channel outputs. The number of output channels can be either 4 (SPLUT-S), 8 (SPLUT-M), or 16 (SPLUT-L). The multi-channel feature maps are then grouped along the channel dimension and fed to either vertical or horizontal aggregation. The aggregated features are then fed to $\rm{LUT}_{HC}$ and $\rm{LUT}_{WC}$, both of which have a $2\times 2$ input pattern. The input pattern of $\rm{LUT}_{HC}$ is along both the height and channel dimensions, while that of $\rm{LUT}_{WC}$ is along both the width and channel dimensions. The aggregation modules, $\rm{LUT}_{WC}$ and $\rm{LUT}_{HC}$, can be repeated multiple times in series to further expand the RF. The final super-resolved (SR) image is produced by summing the outputs of the two branches. The proposed HKLUT architecture is illustrated in Fig. 12 (d). Our method utilizes the two-branch structure of SPLUT to eliminate the interpolation step. However, instead of employing the same set of LUTs in both branches, we propose an asymmetric parallel structure that uses smaller LUTs for the LSB branch to reduce storage space. Within each branch, we use multiple LUTs with the rotation ensemble trick similar to MuLUT. However, LUTs of our design have fewer input pixels, resulting in an exponential reduction in size. Summing the results from the two branches yields the final SR image. The entire structure is treated as a building block. A multistage structure is constructed by repeating the whole module. This allows convenient implementation of different progressive upsample strategies to further reduce storage.
Appendix E Lightweight CNN Comparisons
In Section 4, we selected RRDB (best accuracy) and CARN-M (storage-accuracy balance) to align with previous literature (SRLUT, SPLUT, and MuLUT) using the same set of baselines. We listed the storage, PSNR (on BSDS100), energy cost, and peak memory of two state-of-the-art (SOTA) lightweight CNNs, FMEN and Omni-SR, the best-performing LUT method (MuLUT-X2), and HKLUT-S for generating a $3\times 1280\times 720$ image through 4×4$\times$4 × SR. In Table 7, we demonstrate that HKLUT-S has the least storage, energy, and memory. Despite a 1dB lower PSNR in LUT compared to CNN (intrinsic to LUT schemes due to inherently restricted RF), LUT-based methods consume significantly lower energy (two orders of magnitude lower, proportional to FLOPs) and memory. This makes them ideal for low-end edge devices without specific hardware like GPU or NPU. We emphasize that our work aims to advance the Pareto front of LUT-based methods in terms of the storage-performance tradeoff. In this context, HKLUT achieves unprecedented sub-1MB results (Fig. 1).
Metrics FMEN Omni-SR MuLUT-X2 HKLUT-S
Storage (KB) 769 792 4160 100
PSNR (dB) 27.63 27.71 26.89 26.73
Energy (µJ) 22630 18432 701 117
Memory (MB) 1024 6712 245 93
Table 7: Comparison between LUT-based approaches and lightweight CNNs. The results are obtained by performing 4×4$\times$4 × SR rendering on a $1280\times 720$ RGB image.
Appendix F Additional Ablation Studies
F.1 Ablation on Residual Connection
Table 8 compares the results of the three commonly used upsampling methods (nearest, bilinear, and bicubic). The effects of the three methods are almost indistinguishable. To reduce runtime, we selected the nearest interpolation as our upsampling strategy for residual connections.
@@ F.1 Ablation on Residual Connection
F.2 Ablation on Architecture Design
To further justify our design choice, we compared HKLUT with a baseline approach that simply combines two SOTA methods, namely SPLUT and MuLUT, by applying MuLUT to both MSB and LSB branches of SPLUT. Table 9 shows that such naïve combination not only requires more storage but also yields inferior performance compared to HKLUT-S.
Method Size Set5 Set14 BSDS100
SPLUT+MuLUT 6144 KB 30.19 27.29 26.70
HKLUT-S 100 KB 30.35 27.39 26.73
Table 9: HKLUT-S outperforms a piece-together of SPLUT and MuLUT with a remarkably smaller size for the 4$\times$ SR task.
Figure 12: Comparison of architectural designs. Residual connections are omitted for brevity. (a) SRLUT (b) SPLUT (c) MuLUT (d) HKLUT (Ours)